# Revisiting Functional Derivatives in Multi-object Tracking

### v1

@@ II-A The Base Space
It is assumed that $\mathcal{X}$ is endowed with a measure-theoretic integration concept. Let $(\mathcal{X},\mathfrak{B},\mu)$ be a measure space with $\mathfrak{B}$ being the Borel $\sigma$-algebra and $\mu$ a reference Radon measure, see Supplement E. If, for instance, $\mathcal{X}{=}\mathbb{R}^{d}$, then it is usual to choose $\mu$ as the $d$-dimensional Lebesgue measure111 More precisely, the restriction of the Lebesgue measure to the standard Borel $\sigma$-algebra on $\mathbb{R}^{d}$. . For simplicity, integration w.r.t. the reference Radon measure $\mu$ is denoted as
@@
If $f$ is a probability density on $\mathcal{X}$, the integral (1) is assumed to be one, which is the usual property one would expect if $\mu$ was the Lebesgue measure. If $\mu$ is translation invariant222 Such a measure cannot be defined on any lcscH space; see [37, Chapter 11]. For $\mathcal{X}{=}\mathbb{R}^{d}$, such $\mu$ is only the Lebesgue measure up to a scaling factor. For more general problems encountered in practice such as for $\mathcal{X}{=}\mathbb{R}^{n_{1}}{\uplus}\mathbb{R}^{n_{2}}$, the Lebesgue measure can be generalized straightforwardly. , it is convenient to introduce units of measurement on $\mathcal{X}$. Denote the unit of measurement in $\mathcal{X}$ as $u$, i.e., $u=\mu(U)$ with $U\in\mathfrak{B}$ being a chosen “unit set” of the space $\mathcal{X}$.
@@ II-B Brief Introduction to RFSs
Any realization $X$ can be seen as an element of a hyperspace comprised of all finite subsets of $\mathcal{X}$, denoted with $\mathcal{F}(\mathcal{X})$. Omitting details333 To introduce the probability measure of $\Xi$ correctly, the space $\mathcal{F}(\mathcal{X})$ has to be turned into a probability space, and $C$ should be taken as a Borel measurable subset of $\mathcal{F}(\mathcal{X})$. This can be done by adopting the Mathéron “hit-or-miss” topology on $\mathcal{F}(\mathcal{X})$ and inducing a $\sigma$-algebra. The details can be found, e.g., in [9, 14, 38][15, Appendix F]. not needed for this paper, the probability measure corresponding to $\Xi$ is denoted with
@@ II-D Density Function
where the set derivative denoted with ${\delta}/{\delta X}$ is iterative, and it is given in Supplement B for completeness. If the cardinality of an RFS is bounded444 In [9], the function $p_{\Xi}(X)$ (5) must vanish identically for all sufficiently large cardinalities $|X|$. It follows that there is no “Poisson RFS” under the formulation of [9] unless it is an approximation, see [9, pp. 223, 235]. , it was shown in [9, pp. 156–168], that the density $p_{\Xi}(X)$ (5) exists under appropriate notion of absolute continuity of $\beta(S)$ (3). Nevertheless, these conditions can be extended to the unbounded RFS framework [9, 14, 15] adopted nowadays, see the Licentiate thesis [38]. If both the densities (4) and (5) exist, the density $p_{\Xi}(X)$ (5), when evaluated at the set $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$, is $\mu$-ae equal to the $n$-th order Janossy density [33, Eq. (38)],
@@ II-F Probability Generating Functional
The PGFL $G_{\Xi}:\Gamma(\mathcal{X})\rightarrow\mathbb{C}$ of an RFS $\Xi$ can then be defined as the expectation555 The PGFL can be defined over a narrower class of certain real-valued functions as a special case [12, Sec. 9.4], but the advantages of doing so are not needed in this paper.  [11, Sec. 5.5], [40]
@@ III-B Definition Almost Everywhere
for any function $q$ whose support666 According to [16, p. 1045], the support of function $q$ is defined as $\mathrm{Supp}(q)\triangleq\{\mathbf{x}\in\mathcal{X}:q(\mathbf{x})\neq 0\}$, cf. Eq. (79) from Supplement D. has $\mu$-measure zero, i.e., $q$ vanishes almost everywhere. This construction assumes that the directional derivative of such functional is moreover a Gâteaux derivative, i.e, it is linear and continuous w.r.t. the direction $g$, see Supplement C. With this, the derivative (14) in the direction of the indicator function $g(\mathbf{x})=\mathbf{1}_{S}(\mathbf{x})$ with $S\subset\mathcal{X}$ being measurable, defines a $\sigma$-finite measure as
@@ III-D Definitions Specific to PGFLs via Distributions
Assuming that $\mathcal{X}=\mathbb{R}^{d}$, the key concept employed in [21, Appendix B.4] and [34] is to use the theory of distributions. Consider a family $\{f_{\mathbf{x},\lambda}\}_{\lambda}$ of functions parameterized by777 To ensure that the unit of measurement of a density on $\mathcal{X}$ (now assumed to be $\mathbb{R}^{d}$) is $u^{-1}$, the unit of measurement of $\lambda$ is taken to be $u$. $\lambda>0$, converging to the Dirac delta centered at $\mathbf{x}\in\mathcal{X}$ as
@@
first, note that for a fixed value of $\lambda$ close to zero, it is likely that $f_{\mathbf{x},\lambda}\notin\Gamma(\mathcal{X})$, but there always exists sufficiently small $\epsilon{>}0$ such that $(\epsilon f_{\mathbf{x},\lambda})\in\Gamma(\mathcal{X})$. Next, notice that for a fixed $h$ such that $\sup_{\mathbf{x}\in\mathcal{X}}|h(\mathbf{x})|<1$ and a fixed $f_{\mathbf{x},\lambda}$, the term $G_{\Xi}[h{+}\epsilon f_{\mathbf{x},\lambda}]$ as a function of $\epsilon$ is analytic in some open region of the complex plane that contains the origin [52, Sec. 4]. The ordinary derivative in (26b) thus exists, but both $h$ and the direction $f_{\mathbf{x},\lambda}$ must be fixed. The properties of the functional derivative (26b) w.r.t. $h$ and $f_{\mathbf{x},\lambda}$ were not studied further999 According to [34], the fact that $G_{\Xi}[h+\epsilon f_{\mathbf{x},\lambda}]$ is analytic as a function of $\epsilon$ around the origin of $\mathbb{C}$ implyes that the resulting functional derivative is Fréchet. This statement is valid only for fixed $h$ and $f_{\mathbf{x},\lambda}$, i.e., the ordinary derivative of $G_{\Xi}[h+\epsilon f_{\mathbf{x},\lambda}]$ w.r.t. $\epsilon$ is Fréchet. This statement is, however, missing the point as $\mathbb{C}$ is a simple example of a Banach space and analytic functions are Fréchet differentiable trivially. in [34]. As a result, it is unclear whether (26) under the current assumptions grants the use of the chain rule. If so, note that (26) must be a Chain derivative instead of Fréchet since the space of distributions is not a normed space as it is a topological vector space whose topology is not even metrizable [53, p. 156].
@@ IV-C Analysis of the Proposed Definition
Let $\mathcal{X}$ be lcH topological space, measures $\eta,\nu_{1},\dots,\nu_{m}{\in}M(\mathcal{X})$, number $p_{0}{\in}\mathbb{C}$ and sequence of functions101010 The functions $p_{n}(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ need not be symmetric for the proof. $p_{n}(\mathbf{x}_{1},\dots,\mathbf{x}_{n})\in C_{0}(\mathcal{X}^{n})$ $\forall n{\in}\mathbb{N}$ such that a finite uniform bound $K$ for all finite cardinalities exists, i.e.,
@@
    Note that positive measures $\eta$ no longer form a linear space111111 For instance, $(-1)\cdot\eta$ is not a positive, but rather a signed measure. that is required for the Fréchet (and also chain differentiability) and thus for the chain rule to hold.
@@ IV-D Open Problems
    The proposed directional and functional derivatives depend on the order of the directions in which the PGFM is being differentiated. That is, to avoid misinterpretation, the notation $\{\,\}$” in $\tfrac{\delta}{\delta\{\mathbf{x}_{1},\dots,\mathbf{x}_{m}\}}$ in (33) and (41) shall be taken with care as was done in (42), (43) and (44). A generalization of the Fubini–Tonelli Theorem for complex measures121212 Note that the symmetry of $p_{\Xi}(X)$ plays no role in this issue. For general measures $\nu_{1},\nu_{2}\in M(\mathcal{X})$, it is possible that $\textstyle\int_{\mathcal{X}}\int_{\mathcal{X}}\nu_{1}(\mathrm{d}\mathbf{x}_{1})\nu_{2}(\mathrm{d}\mathbf{x}_{2})\neq\int_{\mathcal{X}}\int_{\mathcal{X}}\nu_{2}(\mathrm{d}\mathbf{x}_{1})\nu_{1}(\mathrm{d}\mathbf{x}_{2})\,.$ (45) , if possible, would help establishing order-independency in (33), (41).
@@ Appendix B Proof of the Proposed Derivative Being Fréchet (Proposition 2)
Let $\beta(S)$ be real, or possibly multivariate (vector-valued) unitless function defined on closed subsets $S\subseteq\mathcal{X}$. To be consistent with the setting of this paper, the original definition of the set derivative [9, pp. 144–150] can be written as131313 In [9], it was assumed that the base space is $\mathcal{X}=\mathbb{R}^{d}\times W$, where $W$ is a finite set. Although the general case of assuming $\mathcal{X}$ to be any lcscH space may appear in the literature [16], the full (i.e., not simplified) definition of set derivatives for this general case is, to the best of our knowledge, unavailable.
@@
This section summarizes the notion of derivatives in abstract spaces, namely the space of functions or measures. Let $\mathcal{Y}$ be a vector (i.e., linear) space and $G:\mathcal{Y}\rightarrow\mathbb{C}$, i.e., a functional. The directional derivative141414 Note that Eq. (73) can be defined more generally, and in the case $G$ is a functional, it is usually called the variational derivative [21, Appendix C.3]. of $G[h]$ in the direction of $g\in\mathcal{Y}$ is given by [54, p. 293]
@@ 2 Topological Vector Spaces: Chain Differential
Generally, there is no notion of gradient in abstract spaces. Historically, the first approach to define gradient for functionals, i.e., functional derivative, could be coined to the Volterra derivative [42, Sec. 2.1]. Note that the calculus of variations [61], [21, Appendix C.3] does not define the functional derivative explicitly as it is not needed, e.g., for solving optimization problems. Assume, on the other hand, there exists a functional transformation $T[h](\mathbf{x})$, using which the directional derivative can be written $\forall g\in\mathcal{Y}$ as151515 In the case $\mathcal{Y}=\mathbb{R}^{d}$, the functional $G[h]$ becomes a multivariate function indexed by $\mathbf{x}$, $g$ becomes a direction vector, and the integral restorts to a sum, thus (77) shrinks to the familiar relation known from undergraduate calculus.
@@
Topological spaces are primarily motivated by the need for describing continuity of functions defined over elements $\mathbf{x}{\in}\mathcal{X}$ of various nature. A topological space $(\mathcal{X},\mathcal{T})$ is a pair, where $\mathcal{X}$ is a nonempty set and $\mathcal{T}$ is a set of all open subsets of $\mathcal{X}$, further called topology161616 Any topology $\mathcal{T}$ satisfies $\emptyset,\mathcal{X}\in\mathcal{T}$ and $\mathcal{T}$ is closed under arbitrary unions and finite intersections of its elements. Unless $\mathcal{X}$ is a singleton, it can be endowed with different topologies.  [37, Chap. 4]. If the topology $\mathcal{T}$ is fixed during the discourse and no confusion may arise, the topological space is denoted only with $\mathcal{X}$ for convenience.
@@ 2 Convenient Sets of Functions on Topological Spaces
Let $(\mathcal{X}_{1},\mathcal{T}_{1})$ and $(\mathcal{X}_{2},\mathcal{T}_{2})$ be two topological spaces. A function $f\colon\mathcal{X}_{1}\to\mathcal{X}_{2}$ is called continuous, if the pre-image171717 By pre-image we mean $f^{-1}(B):=\{x\in\mathcal{X}_{1}\colon f(x)\in B\}$. $f^{-1}(B)$ is from $\mathcal{T}_{1}$ for any set $B\in\mathcal{T}_{2}$ [37, p. 119]. When $\mathcal{X}_{2}$ is $\mathbb{R}$ or $\mathbb{C}$, the topology $\mathcal{T}_{2}$ is assumed to be the standard topology induced by the usual metric on $\mathbb{R}$ or $\mathbb{C}$, i.e., by d(x,y)=|yx|d(x,y)=|y-x|.

### v3

@@ II-A The Base Space
It is assumed that $\mathcal{X}$ is endowed with a measure-theoretic integration concept. Let $(\mathcal{X},\mathfrak{B},\mu)$ be a measure space with $\mathfrak{B}$ being the Borel $\sigma$-algebra and $\mu$ a reference Radon measure, see Supplement E. If, for instance, $\mathcal{X}{=}\mathbb{R}^{d}$, then it is usual to choose $\mu$ as the $d$-dimensional Lebesgue measure [^1] More precisely, the restriction of the Lebesgue measure to the standard Borel $\sigma$-algebra on $\mathbb{R}^{d}$. . For simplicity, integration w.r.t. the reference Radon measure $\mu$ is denoted as
@@
If $f$ is a probability density on $\mathcal{X}$, the integral (1) is assumed to be one, which is the usual property one would expect if $\mu$ was the Lebesgue measure. If $\mu$ is translation invariant [^2] Such a measure cannot be defined on any lcscH space; see [37, Chapter 11]. For $\mathcal{X}{=}\mathbb{R}^{d}$, such $\mu$ is only the Lebesgue measure up to a scaling factor. For more general problems encountered in practice such as for $\mathcal{X}{=}\mathbb{R}^{n_{1}}{\uplus}\mathbb{R}^{n_{2}}$, the Lebesgue measure can be generalized straightforwardly. , it is convenient to introduce units of measurement on $\mathcal{X}$. Denote the unit of measurement in $\mathcal{X}$ as $u$, i.e., $u=\mu(U)$ with $U\in\mathfrak{B}$ being a chosen “unit set” of the space $\mathcal{X}$.
@@ II-B Brief Introduction to RFSs
Any realization $X$ can be seen as an element of a hyperspace comprised of all finite subsets of $\mathcal{X}$, denoted with $\mathcal{F}(\mathcal{X})$. Omitting details [^3] To introduce the probability measure of $\Xi$ correctly, the space $\mathcal{F}(\mathcal{X})$ has to be turned into a probability space, and $C$ should be taken as a Borel measurable subset of $\mathcal{F}(\mathcal{X})$. This can be done by adopting the Mathéron “hit-or-miss” topology on $\mathcal{F}(\mathcal{X})$ and inducing a $\sigma$-algebra. The details can be found, e.g., in [9, 14, 38][15, Appendix F]. not needed for this paper, the probability measure corresponding to $\Xi$ is denoted with
@@ II-D Density Function
where the set derivative denoted with ${\delta}/{\delta X}$ is iterative, and it is given in Supplement B for completeness. If the cardinality of an RFS is bounded [^4] In [9], the function $p_{\Xi}(X)$ (5) must vanish identically for all sufficiently large cardinalities $|X|$. It follows that there is no “Poisson RFS” under the formulation of [9] unless it is an approximation, see [9, pp. 223, 235]. , it was shown in [9, pp. 156–168], that the density $p_{\Xi}(X)$ (5) exists under appropriate notion of absolute continuity of $\beta(S)$ (3). Nevertheless, these conditions can be extended to the unbounded RFS framework [9, 14, 15] adopted nowadays, see the Licentiate thesis [38]. If both the densities (4) and (5) exist, the density $p_{\Xi}(X)$ (5), when evaluated at the set $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$, is $\mu$-ae equal to the $n$-th order Janossy density [33, Eq. (38)],
@@ II-F Probability Generating Functional
The PGFL $G_{\Xi}:\Gamma(\mathcal{X})\rightarrow\mathbb{C}$ of an RFS $\Xi$ can then be defined as the expectation [^5] The PGFL can be defined over a narrower class of certain real-valued functions as a special case [12, Sec. 9.4], but the advantages of doing so are not needed in this paper.  [11, Sec. 5.5], [40]
@@ III-B Definition Almost Everywhere
for any function $q$ whose support [^6] According to [16, p. 1045], the support of function $q$ is defined as $\mathrm{Supp}(q)\triangleq\{\mathbf{x}\in\mathcal{X}:q(\mathbf{x})\neq 0\}$, cf. Eq. (79) from Supplement D. has $\mu$-measure zero, i.e., $q$ vanishes almost everywhere. This construction assumes that the directional derivative of such functional is moreover a Gâteaux derivative, i.e, it is linear and continuous w.r.t. the direction $g$, see Supplement C. With this, the derivative (14) in the direction of the indicator function $g(\mathbf{x})=\mathbf{1}_{S}(\mathbf{x})$ with $S\subset\mathcal{X}$ being measurable, defines a $\sigma$-finite measure as
@@ III-D Definitions Specific to PGFLs via Distributions
Assuming that $\mathcal{X}=\mathbb{R}^{d}$, the key concept employed in [21, Appendix B.4] and [34] is to use the theory of distributions. Consider a family $\{f_{\mathbf{x},\lambda}\}_{\lambda}$ of functions parameterized by [^7] To ensure that the unit of measurement of a density on $\mathcal{X}$ (now assumed to be $\mathbb{R}^{d}$) is $u^{-1}$, the unit of measurement of $\lambda$ is taken to be $u$. $\lambda>0$, converging to the Dirac delta centered at $\mathbf{x}\in\mathcal{X}$ as
@@
first, note that for a fixed value of $\lambda$ close to zero, it is likely that $f_{\mathbf{x},\lambda}\notin\Gamma(\mathcal{X})$, but there always exists sufficiently small $\epsilon{>}0$ such that $(\epsilon f_{\mathbf{x},\lambda})\in\Gamma(\mathcal{X})$. Next, notice that for a fixed $h$ such that $\sup_{\mathbf{x}\in\mathcal{X}}|h(\mathbf{x})|<1$ and a fixed $f_{\mathbf{x},\lambda}$, the term $G_{\Xi}[h{+}\epsilon f_{\mathbf{x},\lambda}]$ as a function of $\epsilon$ is analytic in some open region of the complex plane that contains the origin [52, Sec. 4]. The ordinary derivative in (26b) thus exists, but both $h$ and the direction $f_{\mathbf{x},\lambda}$ must be fixed. The properties of the functional derivative (26b) w.r.t. $h$ and $f_{\mathbf{x},\lambda}$ were not studied further [^9] According to [34], the fact that $G_{\Xi}[h+\epsilon f_{\mathbf{x},\lambda}]$ is analytic as a function of $\epsilon$ around the origin of $\mathbb{C}$ implyes that the resulting functional derivative is Fréchet. This statement is valid only for fixed $h$ and $f_{\mathbf{x},\lambda}$, i.e., the ordinary derivative of $G_{\Xi}[h+\epsilon f_{\mathbf{x},\lambda}]$ w.r.t. $\epsilon$ is Fréchet. This statement is, however, missing the point as $\mathbb{C}$ is a simple example of a Banach space and analytic functions are Fréchet differentiable trivially. in [34]. As a result, it is unclear whether (26) under the current assumptions grants the use of the chain rule. If so, note that (26) must be a Chain derivative instead of Fréchet since the space of distributions is not a normed space as it is a topological vector space whose topology is not even metrizable [53, p. 156].
@@ IV-C Analysis of the Proposed Definition
Let $\mathcal{X}$ be lcH topological space, measures $\eta,\nu_{1},\dots,\nu_{m}{\in}M(\mathcal{X})$, number $p_{0}{\in}\mathbb{C}$ and sequence of functions [^10] The functions $p_{n}(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ need not be symmetric for the proof. $p_{n}(\mathbf{x}_{1},\dots,\mathbf{x}_{n})\in C_{0}(\mathcal{X}^{n})$ $\forall n{\in}\mathbb{N}$ such that a finite uniform bound $K$ for all finite cardinalities exists, i.e.,
@@
    Note that positive measures $\eta$ no longer form a linear space [^11] For instance, $(-1)\cdot\eta$ is not a positive, but rather a signed measure. that is required for the Fréchet (and also chain differentiability) and thus for the chain rule to hold.
@@ IV-D Open Problems
    The proposed directional and functional derivatives depend on the order of the directions in which the PGFM is being differentiated. That is, to avoid misinterpretation, the notation $\{\,\}$” in $\tfrac{\delta}{\delta\{\mathbf{x}_{1},\dots,\mathbf{x}_{m}\}}$ in (33) and (41) shall be taken with care as was done in (42), (43) and (44). A generalization of the Fubini–Tonelli Theorem for complex measures [^12] Note that the symmetry of $p_{\Xi}(X)$ plays no role in this issue. For general measures $\nu_{1},\nu_{2}\in M(\mathcal{X})$, it is possible that $\textstyle\int_{\mathcal{X}}\int_{\mathcal{X}}\nu_{1}(\mathrm{d}\mathbf{x}_{1})\nu_{2}(\mathrm{d}\mathbf{x}_{2})\neq\int_{\mathcal{X}}\int_{\mathcal{X}}\nu_{2}(\mathrm{d}\mathbf{x}_{1})\nu_{1}(\mathrm{d}\mathbf{x}_{2})\,.$ (45) , if possible, would help establishing order-independency in (33), (41).
@@ Appendix B Proof of the Proposed Derivative Being Fréchet (Proposition 2)
Let $\beta(S)$ be real, or possibly multivariate (vector-valued) unitless function defined on closed subsets $S\subseteq\mathcal{X}$. To be consistent with the setting of this paper, the original definition of the set derivative [9, pp. 144–150] can be written as [^13] In [9], it was assumed that the base space is $\mathcal{X}=\mathbb{R}^{d}\times W$, where $W$ is a finite set. Although the general case of assuming $\mathcal{X}$ to be any lcscH space may appear in the literature [16], the full (i.e., not simplified) definition of set derivatives for this general case is, to the best of our knowledge, unavailable.
@@
This section summarizes the notion of derivatives in abstract spaces, namely the space of functions or measures. Let $\mathcal{Y}$ be a vector (i.e., linear) space and $G:\mathcal{Y}\rightarrow\mathbb{C}$, i.e., a functional. The directional derivative [^14] Note that Eq. (73) can be defined more generally, and in the case $G$ is a functional, it is usually called the variational derivative [21, Appendix C.3]. of $G[h]$ in the direction of $g\in\mathcal{Y}$ is given by [54, p. 293]
@@ 2 Topological Vector Spaces: Chain Differential
Generally, there is no notion of gradient in abstract spaces. Historically, the first approach to define gradient for functionals, i.e., functional derivative, could be coined to the Volterra derivative [42, Sec. 2.1]. Note that the calculus of variations [61], [21, Appendix C.3] does not define the functional derivative explicitly as it is not needed, e.g., for solving optimization problems. Assume, on the other hand, there exists a functional transformation $T[h](\mathbf{x})$, using which the directional derivative can be written $\forall g\in\mathcal{Y}$ as [^15] In the case $\mathcal{Y}=\mathbb{R}^{d}$, the functional $G[h]$ becomes a multivariate function indexed by $\mathbf{x}$, $g$ becomes a direction vector, and the integral restorts to a sum, thus (77) shrinks to the familiar relation known from undergraduate calculus.
@@
Topological spaces are primarily motivated by the need for describing continuity of functions defined over elements $\mathbf{x}{\in}\mathcal{X}$ of various nature. A topological space $(\mathcal{X},\mathcal{T})$ is a pair, where $\mathcal{X}$ is a nonempty set and $\mathcal{T}$ is a set of all open subsets of $\mathcal{X}$, further called topology [^16] Any topology $\mathcal{T}$ satisfies $\emptyset,\mathcal{X}\in\mathcal{T}$ and $\mathcal{T}$ is closed under arbitrary unions and finite intersections of its elements. Unless $\mathcal{X}$ is a singleton, it can be endowed with different topologies.  [37, Chap. 4]. If the topology $\mathcal{T}$ is fixed during the discourse and no confusion may arise, the topological space is denoted only with $\mathcal{X}$ for convenience.
@@ 2 Convenient Sets of Functions on Topological Spaces
Let $(\mathcal{X}_{1},\mathcal{T}_{1})$ and $(\mathcal{X}_{2},\mathcal{T}_{2})$ be two topological spaces. A function $f\colon\mathcal{X}_{1}\to\mathcal{X}_{2}$ is called continuous, if the pre-image [^17] By pre-image we mean $f^{-1}(B):=\{x\in\mathcal{X}_{1}\colon f(x)\in B\}$. $f^{-1}(B)$ is from $\mathcal{T}_{1}$ for any set $B\in\mathcal{T}_{2}$ [37, p. 119]. When $\mathcal{X}_{2}$ is $\mathbb{R}$ or $\mathbb{C}$, the topology $\mathcal{T}_{2}$ is assumed to be the standard topology induced by the usual metric on $\mathbb{R}$ or $\mathbb{C}$, i.e., by d(x,y)=|yx|d(x,y)=|y-x|.